OpenStudy (anonymous):

I need to check if the vertex is right-- y>-3|-x-4|+4 Is the vertex (4,-4) or am I way off? I also need to graph this (preferably using an x y chart so if it's alright could someone also please clarify how exactly you use one? Thanks!

jimthompson5910 (jim_thompson5910):

y = -3|-x-4|+4 y = -3|-1(x+4)|+4 y = -3|-1|*|x+4|+4 y = -3*1*|x+4|+4 y = -3|x+4|+4 y = -3|x-(-4)|+4 This equation is now in the form y = a|x-h| + k where the vertex of this general equation is (h,k) Since in this case, h = -4 and k = 4, the vertex is (-4, 4)

jimthompson5910 (jim_thompson5910):

Because a = -3, which is negative, we know that this V shaped graph will open downward

jimthompson5910 (jim_thompson5910):

To graph y = -3|-x-4|+4, you need a series of points To get any point, you plug in a number for x to get a corresponding value for y example when x = 0, y is... y = -3|-x-4|+4 y = -3|-0-4|+4 ... replace x with 0 (since x = 0) y = -3|0-4| + 4 y = -3|-4| + 4 y = -3(4) + 4 y = -12 + 4 y = -8 meaning that the point (0,-8) is on the graph

jimthompson5910 (jim_thompson5910):

you repeat this for other x values to get other points (the more points, the more accurate the graph is)

OpenStudy (anonymous):

Alright, thanks so much!

jimthompson5910 (jim_thompson5910):

you're welcome